# Redistribution Through

### v2

@@ Appendix C Proof of Proposition 2
$\Leftarrow$ Suppose that $\sigma\in\Sigma_{\Omega}(\mu)$ satisfies points (i) and (ii) in Proposition 2. We first show that there exists no $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ that can be obtained from $\sigma$ by a redistributive buyer transfer. By contradiction, this would imply a segmentation $\sigma^{\prime}$ obtained from $\sigma$ by a transfer of an $\varepsilon$ mass from $(\theta^{\prime},p^{\prime\prime})$ and $(\theta^{\prime\prime},p^{\prime})$ with $p^{\prime}<p^{\prime\prime}$ and $\theta^{\prime}<\theta^{\prime\prime}$ to $(\theta^{\prime},p^{\prime})$ and $(\theta^{\prime\prime},p^{\prime\prime})\in\operatorname{supp}(\sigma)$. Claim (ii) in Proposition 2 implies that $\theta^{\prime\prime}\in\mathcal{I}_{\sigma}(p^{\prime\prime})$, which in turn means that such a $\sigma^{\prime}$ cannot satisfy Ob.
We now show that there exists no $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ that can be obtained from $\sigma$ by a downward buyer transfer. By contradiction take $(\theta^{\prime},p^{\prime\prime})\in\operatorname{supp}(\sigma)$, $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p^{\prime}<p^{\prime\prime}$, $\varepsilon>0$, and suppose that $\sigma^{\prime}$ obtained from $\sigma$ by a downward buyer transfer from $(\theta^{\prime},p^{\prime\prime})$ to $(\theta^{\prime},p^{\prime})$ is in $\Sigma_{\Omega}(\mu)$. Claim (i) in Proposition 2 implies that there exists $q>p^{\prime}$ such that $q\in\mathcal{I}_{\sigma}(p^{\prime})$. If $q\leq\theta^{\prime}$, then we must have
@@
which contradicts $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$. Therefore, it must be that $q>\theta^{\prime}\geq p^{\prime\prime}$. Now, claim (ii) in Proposition 2 implies that $q\in\mathcal{I}_{\sigma}(p^{\prime\prime})$, meaning that
@@
which contradicts $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$. ∎
Appendix D Proof of Theorem 2.B
Proof of claim (ii).
$\Leftarrow$ Consider a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ that is $\succsim_{\mathcal{R}}$-maximal but not strongly monotone. Thus, there exists at least two segments $p,\theta_{i}\in\operatorname{supp}(\sigma_{P})$ such that $\max\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}>\theta_{i}$. We denote as $\theta^{\prime}$ the lowest type in the segment $p$ that is strictly above $\theta$, and as $\theta_{k}$ the maximal type in the segment $\theta_{i}$:
@@
where both strict inequalities come from $w\in\overline{\mathcal{R}}$. In addition, given that $\sigma$ is $\succsim_{\mathcal{R}}$-maximal, $\sigma+r$ does not satisfy Ob. In particular, there exists $\hat{\theta}$, with $\theta_{i}<\hat{\theta}\leq\theta^{\prime}$, such that $\hat{\theta}\in\mathcal{I}^{\sigma}(\theta_{i})$ (see Appendix C for the definition of $\mathcal{I}^{\sigma}(\theta)$), and thus:
@@
where the first inequality follows from $\theta_{i}$ being an optimal price in segment $\theta_{i}$ according $\sigma$, and the second inequality follows $q\geq\theta_{i+1}$.
Therefore, $\sigma+r+t\in\Sigma_{\Omega}(\mu)$. The loss from the transfer $t$ is
@@
Therefore, implementing the transfers $r$ and $t$ is beneficial if and only if 9 $\geq$ 11. If $w$ is strongly redistributive, we have that 10 $\geq$ 11. Given that 9 $\geq$ 10, this ends the proof.
$\Rightarrow$ We prove the contrapositive. Suppose that $w\in\overline{\mathcal{R}}$ is not strongly redistributive. So there exist $p<\theta_{i}<\theta_{i+1}\leq\theta_{k}$, such that
$\bigl{[}w(\theta_{i},p)-w(\theta_{i},\theta_{i})\bigr{]}-\bigl{[}w(\theta_{i+1},p)-w(\theta_{i+1},\theta_{i})\bigr{]}<\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\bigl{[}w(\theta_{k},\theta_{i})-w(\theta_{k},\theta_{k})\bigr{]}.$ (12)
Let $\mu\in\Delta\bigl{(}\{p,\theta_{i},\theta_{i+1},\theta_{k}\}\bigr{)}$ such that the unique strongly monotone and $\succsim_{\mathcal{R}}$-maximal segmentation, denoted $\sigma$, has three segments:
@@
It is easy to show that such a $\mu$ exists. [^30] It can be provided upon requests. Let $\varepsilon>0$ be sufficiently small. We will now perform two transfers. The first one is a downward transfer $d$ defined by:
@@
$s(\theta_{i},\theta_{i})=s(\theta_{i+1},p)=-s(\theta_{i+1},\theta_{i})=-s(\theta_{i},p)-\varepsilon,$
@@
$\theta_{i}\left(\sigma(\theta_{i},\theta_{i})+\sigma(\theta_{i},\theta_{i+1})+\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)$ $-\theta_{i+1}\left(\sigma(\theta_{i},\theta_{i+1})-\varepsilon+\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)$
@@
where the first equality follows from $\theta_{i+1}$ being an optimal price in segment $\theta_{i}$ for segmentation $\sigma$. Similarly, one can show that charging $\theta_{i}$ always gives a greater profit than charging $\theta_{k}$ if the segmentation is $\sigma+d+s$. Therefore $\sigma+d+s\in\Sigma_{\Omega}(\mu)$. Furthermore, it directly follows from Equation 12 that $\langle w,\sigma+d+s\rangle>\langle w,\sigma\rangle$. Therefore, $\sigma$, which is the unique strongly monotone and $\succsim_{\mathcal{R}}$-maximal segmentation of $\mu$ (Lemma 2) is not optimal for $w$. ∎
Appendix E Proof of Lemma 2
@@
First, if the uniform price of $\mu$ is $\theta_{1}$, there exists unique segmentation that is $\succsim_{\mathcal{R}}$-maximal, the one that does not segment the market. Suppose now that the uniform price of $\mu$ is strictly greater than $\theta_{1}$. Let $\sigma\in\Sigma_{\Omega}(\mu)$ be strongly monotone and $\succsim_{\mathcal{R}}$-maximal. Therefore it satisfies the conditions of Proposition 2. Denote as $\theta_{k}$ the maximal type that is in the segment $\theta_{1}$. By Proposition 2, there exists $\theta_{i}$, with $\theta_{1}<\theta_{i}\leq\theta_{k}$ such that $\theta_{i}\in\mathcal{I}_{\sigma}(\theta_{1})$. Consequently, any segmentation that is strongly monotone and $\succsim_{\mathcal{R}}$-maximal must have the same market segment $\theta_{1}$. By iterating the same argument for the other segments (starting with the next segment $\theta_{k}$ or $\theta_{k+1}$), this implies that $\sigma$ is the unique segmentation that satisfies these conditions. ∎
Appendix F Pseudo-code for the greedy redistributive algorithm
Algorithm 1 Greedy redistributive algorithm
$\sigma\leftarrow 0_{K\times K}$; $\triangleright$ Initiates the segmentation as being the null matrix of size $K\times K$. $\sigma_{i,j}$ denotes the element of $\sigma$ in row $i$ and column $j$ and $\sigma_{i,\bullet}$ denotes the row $i$ of $\sigma$ seen as a vector in $\mathbb{R}^{K}$.
[…]
$\delta_{k}\leftarrow\delta_{k}+x$;
[…]
$\sigma_{p,t}\leftarrow m$; $\triangleright$ Allocate this mass to segment $p$.
$p\leftarrow t$; $\triangleright$ Initiate a new segment with price $t$.
[…]

### v5

@@ Appendix C Proof of Lemma˜3
[…]
We now show that there exists no $\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$ that can be obtained from $\sigma$ by a downward buyer transfer. By contradiction take $(\theta^{\prime},p^{\prime\prime})\in\operatorname{supp}(\sigma)$, $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p^{\prime}<p^{\prime\prime}$, $\varepsilon>0$, and suppose that $\sigma^{\prime}$ obtained from $\sigma$ by a downward buyer transfer from $(\theta^{\prime},p^{\prime\prime})$ to $(\theta^{\prime},p^{\prime})$ is in $\Sigma^{\mathrm{eff}}(\mu)$. Part (a) in Definition˜1 implies that there exists $q>p^{\prime}$ such that $q\in\mathcal{I}_{\sigma}(p^{\prime})$. If $q\leq\theta^{\prime}$, then we must have
@@
which contradicts $\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$. Therefore, it must be that $q>\theta^{\prime}\geq p^{\prime\prime}$. Now, part (b) in Definition˜1 implies that $q\in\mathcal{I}_{\sigma}(p^{\prime\prime})$, meaning that
@@
which contradicts $\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$. ∎
Appendix D Proof of Proposition˜2
Proof.
[…]
@@ Appendix D Proof of Proposition˜2
[…]
@@
where the first inequality follows from $\theta_{i}$ being an optimal price in segment $\theta_{i}$ according to $\sigma$, and the second inequality follows from $q\geq\theta_{i+1}$.
Therefore, $\sigma+r+t\in\Sigma^{\mathrm{eff}}(\mu)$. The loss from the transfer $t$ is
@@
Therefore, implementing the transfers $r$ and $t$ is beneficial if and only if 6 $\geq$ 8. If $w$ is strongly redistributive, we have that 7 $\geq$ 8. Given that 6 $\geq$ 7, this ends the proof.
[…]
$\bigl[w(\theta_{i},p)-w(\theta_{i},\theta_{i})\bigr]$- $\bigl[w(\theta_{i+1},p)-w(\theta_{i+1},\theta_{i})\bigr]$ (9)
$<\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\bigl[w(\theta_{k},\theta_{i})-w(\theta_{k},\theta_{k})\bigr].$
[…]
@@
Such a $\mu$ exists. For instance, let $\mu(p)>0$ and $\mu(\theta_{i+1})>0$ be arbitrarily small, and set $\mu(\theta_{i})=\frac{p}{\theta_{i}-p}\mu(p)+\frac{\theta_{i+1}-\theta_{i}}{\theta_{i}}\mu(\theta_{i+1})$ and $\mu(\theta_{k})=1-\mu(p)-\mu(\theta_{i})-\mu(\theta_{i+1})$. One can verify that this $\mu$ satisfies all the desired conditions. Let $\varepsilon>0$ be sufficiently small. We will now perform two transfers. The first one is a downward transfer $d$ defined by:
@@
$s(\theta_{i},\theta_{i})=s(\theta_{i+1},p)=-s(\theta_{i+1},\theta_{i})=-s(\theta_{i},p)=\varepsilon,$
@@
$\theta_{i}\left(\sigma(\theta_{i},\theta_{i})+\sigma(\theta_{i+1},\theta_{i})+\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)$ $-\theta_{i+1}\left(\sigma(\theta_{i+1},\theta_{i})-\varepsilon+\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)$
@@
[…]
Appendix E Proof of Lemma˜4
@@
First, if the uniform price of $\mu$ is $\theta_{1}$, there exists a unique segmentation that is saturated, the one that does not segment the market. Suppose now that the uniform price of $\mu$ is strictly greater than $\theta_{1}$. Let $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ be strongly monotone and saturated. Denote as $\theta_{k}$ the maximal type that is in the segment $\theta_{1}$. By Definition˜1, there exists $\theta_{i}$, with $\theta_{1}<\theta_{i}\leq\theta_{k}$ such that $\theta_{i}\in\mathcal{I}_{\sigma}(\theta_{1})$. Consequently, any segmentation that is strongly monotone and saturated must have the same market segment $\theta_{1}$. Conclude by iterating the same argument for the other segments. ∎